OpenStudy (anonymous):

Find an expression for f^-1 (x) (inverse) y = (2x + 7) / (x + 2)

OpenStudy (turingtest):

I'm thinking you didn't see how to factor out the x\[y=\frac{2x+7}{x+2}\]\[yx+2y=2x+7\]\[yx-2x=7-2y\]\[x(y-2)=7-2y\]pretty sure you got it now...

OpenStudy (anonymous):

You were right :p I couldn't factor it out... I feel so stupid! Thanks! :)

OpenStudy (turingtest):

It happens to us all. Now you know though, so next time no excuse! :P